# Segre forms of singular metrics on vector bundles and Lelong numbers

### v1

@@ Abstract.
Let $E\to X$ be a holomorphic vector bundle with a singular Griffiths negative Hermitian metric $h$ with analytic singularities. One can define, given a smooth reference metric $h_{0}$, a current $s(E,h,h_{0})$ that is called the associated Segre form, which defines the expected Bott-Chern class and coincides with the usual Segre form of $h$ where it is smooth. We prove that $s(E,h,h_{0})$ is the limit of the Segre forms of a sequence of smooth metrics if the metric is smooth outside the unbounded locus, and in general as a limit of Segre forms of metrics with empty unbounded locus. One can also define an associated Chern form $c(E,h,h_{0})$ and we prove that the Lelong numbers of $s(E,h,h_{0})$ and $c(E,h,h_{0})$ are integers if the singularities are integral, and non-negative for $s(E,h,h_{0})$. We also extend these results to a slightly larger class of singular metrics with analytic singularities.
@@ 1. Introduction
Inspired by the importance during several decades of singular metrics on line bundles in algebraic geometry, with origins in [12], several efforts have been done to study singular metrics on higher rank holomorphic vector bundles. In particular, such metrics appeared in the work of Berndtsson-Păun, [9], in connection with pseudoeffectivity of twisted relative canonical bundles. Contrary to the line bundle case, it is not clear what the curvature and connection matrix for a singular metric should be, cf. [20, Theorem 1.5]. This would appear to be an obstacle for defining associated Chern forms (or more general characteristic forms like, e.g., Segre forms), since the standard definition is as certain polynomials in the entries of the curvature matrix. However, for appropriate metrics, it has still been possible to introduce such associated forms, or more precisely currents of order 00, see, e.g., [15],[16],[17],[18] and [21]. In this article, we will further the study of such associated forms in more general situations.
Assume that $E\to X$ is a holomorphic vector bundle over a complex manifold $X$ with a Griffiths negative singular metric $h$ with analytic singularities; this means that $\log|\cdot|_{h}$ is plurisubharmonic (psh) with analytic singularities on $E\setminus 0_{E}$, see Section 3. In [18] was introduced an associated Segre form [^1] We keep the notion ’Segre form’ although in the singular case $s(E,h,h_{0})$ is in general a current (of order 00). that represents the expected Bott-Chern class, and coincides with the usual Segre form where the metric is non-singular. The form $s(E,h,h_{0})$ slightly depends on a choice of smooth reference metric $h_{0}$ on $E$. For a singular metric $h$, we let the unbounded locus $Z=Z_{h}\subset X$ denote the complement of the set of $x\in X$ where $\log|\cdot|^{2}_{h}$ is locally bounded on $E\setminus 0_{E}$ above a neighborhood of $x$. Assume for the moment that in addition $h$ is smooth outside of $Z$. Let $h_{\epsilon}=h+{\epsilon}h_{0}$, i.e.,
@@
Then $h_{\epsilon}$ is a smooth metric on $E$. Thus the associated Segre forms $s(E,h_{\epsilon})$ are well-defined and our first main result states that
@@
This further suggests that the definition of $s(E,h,h_{0})$ is natural.
In this paper we will consider bundles with metrics such that $\psi=\log|\cdot|^{2}$ is quasi-plurisubharmonic (qpsh) with analytic singularities on $E$. The latter condition means that locally $\psi=c\log|f|^{2}+b$, where $f$ is a tuple of holomorphic functions, $c>0$ and $b$ is bounded. We say that the singularities are integral (rational/real) if $c$ is a positive integer (rational/real number). Such a metric, as well as the associated bundle will be called qas (qpsh with analytic singularities). For such metrics, the unbounded locus will be an analytic subset of $X$. We extend the definition of Segre form $s(E,h,h_{0})$ to qas bundles $E\to X$. For simplicity, we will assume that qas metrics have integral singularities, which is practical in order to keep track on arithmetic information. However, almost all the results extend easily to the case of rational or real singularities, see Remark 6.9.
@@
Assume that $g\colon E\to F$ is a holomorphic vector bundle morphism and $F$ has a smooth metric, or a qas metric with empty unbounded locus. Then the metric $|\alpha|^{2}=|g(x)\alpha|^{2}$ on $E$ inherited from $F$ is qas, see Corollary 3.4. ∎
@@
An example of a singular metric $e^{\psi}$ on a line bundle $L$ is given by $\psi=\log(|\sigma_{1}|^{2}+\cdots+|\sigma_{m}|^{2})$, where $\sigma_{j}$ are sections of $L^{-1}$. It is covered by Example 1.1 via the mapping $L\to X\times\oplus_{1}^{m}{\mathbb{C}}$, defined by $(\sigma_{1},\ldots,\sigma_{m})$. Such metrics are important in modern complex algebraic geometry. ∎
By a smooth metric $h$ on $E$, we mean (as usual) that with respect to a local trivialization, the matrix representing $h$ has entries which are smooth functions, and $h$ is strictly positive definite. We begin by considering the basic case of qas metrics with empty unbounded locus, i.e., which are still strictly positive definite, bot not necessarily smooth. Then we have the following result.
Proposition 1.3.
Assume that $E\to X$ is a holomorphic vector bundle equipped with a qas metric $h$, and assume that the unbounded locus $Z$ of $h$ is empty. Then the Segre form $s(E,h,h_{0})$ associated to $h$ is independent of the choice of smooth reference metric $h_{0}$, and may hence be denoted $s(E,h)$. Furthermore, there exists locally a sequence of smooth metrics $h_{\epsilon}$ decreasing to $h$ such that
(1.3) $\lim_{{\epsilon}\to 0}s(E,h_{\epsilon})=s(E,h).$
If $h$ is additionally Griffiths negative, then this is a special case of the results in [17].
We consider now a general qas metric $h$ on $E$. Then, the metric $h_{\epsilon}$ defined by (1.1) is a qas metric with empty unbounded locus, cf. Lemma 3.5 (iv). Hence, $s(E,h_{\epsilon})$ is well-defined by the above proposition, independent of any choice of smooth reference metric.
The components $s_{k}(E,h,h_{0})$ of the Segre form are closed currents of order 00, but not necessarily positive unless both $h$ and $h_{0}$ are Griffiths negative. There are anyway well-defined multiplicities $\operatorname{mult}_{x}s_{k}(E,h,h_{0})$ at each point $x$, defined in the same way as Lelong numbers, see Section 4, and they turn out to be non-negative integers. Here is our first main theorem:
@@
Assume that $E\to X$ is a holomorphic vector bundle equipped with a qas metric $h$, and let $Z$ be the unbounded locus of $h$. Let $h_{0}$ be a smooth reference metric, and let $s(E,h,h_{0})$ denote the associated Segre form.
(i) If $h_{\epsilon}$ is the metric defined by (1.1), then
$s(E,h_{\epsilon})\to s(E,h,h_{0}).$
@@
where $M^{E,h,h_{0}}$ has support on $Z$ and $s^{\prime}(E,h)={\bf 1}_{X\setminus Z}s(E,h,h_{0})$ is independent of $h_{0}$ and if $h$ is smooth on $X\setminus Z$, $s^{\prime}(E,h)$ coincides with the usual Segre form of $h$ there.
@@
(vi) For each integer $k\geq 0$, there is a decomposition
@@
where $S_{k}^{E,h}$ is the Lelong current of an analytic cycle of codimension $k$, and the non-vanishing multiplicities of $s_{k}^{\prime}$ and $N_{k}$ are contained in an analytic set with codimension $\geq k+1$. The multiplicities of $s_{k}^{\prime}$, $S_{k}$ and $N_{k}$ are non-negative integers, independent of the choice of $h_{0}$.
(vii) If $h^{\prime}$ is a metric on $E\to X$ comparable to $h$, and $h_{0}^{\prime}$ is a smooth metric on $E$, then $\operatorname{mult}_{x}M_{k}^{E,h,h_{0}}=\operatorname{mult}_{x}M_{k}^{E,h^{\prime},h_{0}^{\prime}}$ for all $x\in X$ and $k\geq 0$, and similarly if $M_{k}$ is replaced by $s_{k}^{\prime}$. In particular $S_{k}^{E,h}=S_{k}^{E,h^{\prime}}$, $k=0,1,2,\ldots$.
@@
(ii) The properties (ii) to (iv) are (at least implicitly) already proved in [18].
@@
(iv) Notice that if $M_{k}^{E,h,h_{0}}$ is a positive current, (1.8) is precisely the Siu decomposition. ∎
In [9], there is a general definition of singular metrics of vector bundles, and a definition of Griffiths negative metrics as mentioned above, and a definition of Griffiths positive metrics, which by definition are dual metrics of Griffiths negative metrics (which coincides with the standard definition when the metric is smooth). In e.g. the references mentioned above concerning definitions of Chern and Segre forms of singular metrics, the results have mainly been formulated for Griffiths positive metrics, which is natural as in the smooth case, Griffiths positive metrics seem to be more frequently of interest than Griffiths negative metrics. In this article, we consider metrics which are generalizations of Griffiths negative metrics instead of Griffiths positive metrics. However, by considering duals of such metrics, one obtains corresponding statements concerning Griffiths positive metrics, cf. Remark 6.2. We opted here for the formulation in the negatively curved case, as it leads to simpler formulations, in particular with less signs and dual metrics.
Recall that in algebraic geometry, characteristic classes of vector bundles such as the Chern class, are defined as endomorphisms on Chow groups, i.e., the ${\mathbb{Z}}$-module of analytic cycles ${\mathcal{Z}}(X)$ modulo rational equivalence. In some cases, as for instance in case of a Hermitian holomorphic vector bundle, there is a canonical representative of its Segre and Chern classes.
Inspired by [5] we introduce a ${\mathbb{Z}}$-module ${\mathcal{QZ}}(X)$ of currents of order 00 that contains the module ${\mathcal{Z}}(X)$ of analytic cycles as a submodule, if cycles are identified with their associated Lelong currents. In particular the $(0,0)$-current $1$ is an element in ${\mathcal{QZ}}(X)$. Any element ${\mathcal{QZ}}(X)$ of a fixed dimension has a well-defined multiplicity (Lelong number) at each point. We also introduce certain quotient classes ${\mathcal{QB}}(X)$ that, contrary to the Chow classes, keep local information such as multiplicities; this fact is crucial in this paper.
A second main result is that the Segre form $s(E,h,h_{0})$ extends to an endomorphism ${\bf s}(E,h,h_{0})$ on ${\mathcal{QZ}}(X)$, such that ${\bf s}_{0}(E,h,h_{0})$ is the identity and ${\bf s}(E,h,h_{0})1=s(E,h,h_{0})$. An important fact is that ${\bf s}(E,h,h_{0})$ defines a well-defined endomorphism ${\bf s}(E,h)$ of ${\mathcal{QB}}(X)$, independent of $h_{0}$. There is an analogue of Theorem 1.4 for ${\bf s}(E,h,h_{0})$, see Theorem 6.5.
Since ${\bf s}(E,h,h_{0})-I$ has positive degree,
@@
is a well-defined endomorphism, such that
It is natural to say that $c(E,h,h_{0}):={\bf c}(E,h,h_{0})1$ is the Chern form associated with $(E,h,h_{0})$. The same definition of $c(E,h,h_{0})$ appeared in [18] but formulated somewhat differently.
@@
Let $(E,h,h_{0})$ be as in Theorem 1.4. Properties (ii) to (vii) of Theorem 1.4 hold if $s(E,h,h_{0})$ is replaced by $c(E,h,h_{0})$, except for that the multiplicities are not necessarily non-negative.
As already mentioned, the starting point for this paper is [18] and the ideas there. The regularization (1.1) relies on results in [6, 10]. The arithmetic analysis of the Lelong numbers relies on ideas in [1], and the modules ${\mathcal{QZ}}(X)$ and ${\mathcal{QB}}(X)$ that are modelled on the somewhat smaller module ${\mathcal{GZ}}(X)$ and ${\mathcal{B}}(X)$ in [5].
@@ Acknowledgements
We are grateful to Håkan Samuelsson Kalm, Antonio Trusiani and David Witt Nyström for helpful discussions.
@@ 2. Preliminaries
Throughout the article, $X$ is assumed to be a complex manifold. Given a holomorphic vector bundle $E\to X$ we let $p\colon{\mathbb{P}}(E)\to X$ denote the projectivization, so that the fiber above $x$ is the set of all lines [^2] There is another common convention in the literature, where the fiber instead is all hyperplanes through the origin. through the origin in $E_{x}$. Then the pullback $p^{*}E\to{\mathbb{P}}(E)$ has the tautological sub(line)bundle $L_{E}\to{\mathbb{P}}(E)$ whose fiber over $(x,[\alpha])\in{\mathbb{P}}(E)$ is the line in $E_{x}$ determined by $\alpha$.
If $E$ has a smooth Hermitian metric $h$, then $p^{*}E$ and hence $L_{E}$ inherit smooth Hermitian metrics. Notice that a local section $s$ of $L_{E}$ has the form $s=\tilde{s}(x,\alpha)\alpha$, where $\tilde{s}$ is a holomorphic function on $E\setminus 0_{E}$ which is $-1$-homogeneous in $\alpha$. Thus any local section can be identified with such an $\tilde{s}$. The induced norm is
@@
Note that even though $\log|\alpha|_{h}^{2}$ is not well-defined on ${\mathbb{P}}(E)$, $dd^{c}\log|\alpha|_{h}^{2}$ is a well-defined form on ${\mathbb{P}}(E)$, and if $s$ is a non-vanishing section of $L_{E}$, then $dd^{c}\log|s|_{h}^{2}=dd^{c}\log|\alpha|_{h}^{2}$. Thus, if $c_{1}(L_{E},e^{\psi})$ and $s_{1}(L_{E},e^{\psi})$ are the first Chern form and Segre form, respectively, then
$-c_{1}(L_{E},e^{\psi})=s_{1}(L_{E},e^{\psi})=dd^{c}\log|\alpha|^{2}_{h},$
so that
$s(L_{E},e^{\psi})=\frac{1}{c(L_{E},e^{\psi})}=\frac{1}{1+c_{1}(L_{E},e^{\psi})}=\frac{1}{1-s_{1}(L_{E},e^{\psi})}=\sum_{\ell=0}^{\infty}(dd^{c}\log|\alpha|_{h}^{2})^{\ell}.$
@@
and it is a smooth form since $p$ is a submersion. We write $s(E,h)=\sum s_{k}(E,h)$, where $s_{k}(E,h)$ has bidegree $(k,k)$. It turns out that $s_{0}(E,h)=1$ and thus $s(E,h)-1$ has positive bidegree. We have the Chern form
@@
By [19] (see also i.e., [14]), these definitions of $c(E,h)$ and $s(E,h)$ coincide with the more standard definitions of Chern form defined as the sum of the elementary symmetric polynomials acting on ($i/2\pi$ times) the curvature tensor $\Theta(h)=D^{2}$ of $(E,h)$, where $D$ denotes the Chern connection with respect to $h$, and the Segre form defined as $s(E,h)=1/c(E,h)$.

### v2

@@ Abstract.
Let $E\to X$ be a holomorphic vector bundle. We consider a class of a singular Hermitian metrics on $E$ with analytic singularities that contains all Griffiths negative such metrics. One can define, given a smooth reference metric $h_{0}$, a current $s(E,h,h_{0})$ called the associated Segre form, which defines the expected Bott-Chern class and coincides with the usual Segre form of $h$ where it is smooth. We prove that $s(E,h,h_{0})$ is the limit of the Segre forms of a sequence of smooth metrics if the metric is smooth outside the degeneracy locus, and in general as a limit of Segre forms of metrics with empty degeneracy locus. One can also define an associated Chern form $c(E,h,h_{0})$. We prove that the Lelong numbers of $s(E,h,h_{0})$ and $c(E,h,h_{0})$ are integers if the singularities are integral, and non-negative for $s(E,h,h_{0})$.
@@ 1. Introduction
Inspired by the importance, during several decades, of singular metrics on line bundles in algebraic geometry, with origins in [12], several efforts have been done to study singular metrics on higher rank holomorphic vector bundles. In particular, such metrics appeared in the work of Berndtsson-Păun, [9], in connection with pseudoeffectivity of twisted relative canonical bundles. Contrary to the line bundle case, it is not clear what the curvature and connection matrix for a singular metric should be, cf. [20, Theorem 1.5]. This would appear to be an obstacle for defining associated Chern forms (or other characteristic forms, like Segre forms), since a standard definition is in terms of the curvature matrix. However, for appropriate metrics, it has still been possible to introduce such associated forms, or more precisely currents of order 0, see, e.g., [15], [16], [17], [18] and [21]. In this article, we further the study of such associated forms, and in slightly more general situations.
Assume that $E\to X$ is a holomorphic vector bundle over a complex manifold $X$ with a Griffiths negative singular metric $h$ with analytic singularities; this means that $\psi=\log|\cdot|^{2}_{h}$ is plurisubharmonic (psh) with analytic singularities on $E\setminus 0_{E}$, see Section 3. The latter condition means that locally $\psi=c\log|f|^{2}+b,$ where $f$ is a tuple of holomorphic functions, $c>0$ and $b$ is bounded. We say that the singularities are integral (rational/real) if $c$ is a positive integer (rational/real number). In [18] was introduced an associated Segre form [^1] We keep the notion ’Segre form’ although in the singular case $s(E,h,h_{0})$ is in general a current (of order 0). $s(E,h,h_{0})$, slightly depending on a smooth reference metric $h_{0}$ on $E$, that represents the expected Bott-Chern class, and coincides with the usual Segre form where the metric is smooth.
We will allow bundles with metrics such that $\psi=\log|\cdot|^{2}_{h}$ is quasi-plurisubharmonic (qpsh) with analytic singularities on $E$. If in addition $\psi$ has integral singularities, we say that the metric, and the bundle equipped with such a metric, is qas. If the singularities are rational or real we say that the bundle/metric is $\mathbb{Q}$}-qas or $\mathbb{R}$}-qas. We mainly consider (integral) qas metrics, which is practical in order to keep track on arithmetic information. There are analogous results for $\mathbb{Q}$}-qas and $\mathbb{R}$}-qas metrics, see Remark 6.10.
The definition of Segre form $s(E,h,h_{0})$ in [18] extends to the case when $h$ is a qas metric on $E\to X$ and $h_{0}$ is a smooth reference metric. Let $h_{\epsilon}=h+{\epsilon}h_{0}$, i.e.,
@@
If $h$ is smooth outside the degeneracy locus, then $h_{\epsilon}$ are smooth and so the classical associated Segre forms $s(E,h_{\epsilon})$ are well-defined. In general $h_{\epsilon}$ is qas with empty degeneracy locus, and the same definition makes sense using Bedford-Taylor products, see (6.18). Our first main result, Theorem 1.3, in particular states that
@@
Assume that $g\colon E\to F$ is a holomorphic vector bundle morphism and $F$ has a smooth metric, or a qas metric with empty degeneracy locus. Then the metric $|\alpha|^{2}=|g(x)\alpha|^{2}$ on $E$ inherited from $F$ is qas, see Corollary 3.6. ∎
@@
An example of a singular metric $e^{\psi}$ on a line bundle $L$ is given by $\psi=\log(|\sigma_{1}|^{2}+\cdots+|\sigma_{m}|^{2})$, where $\sigma_{j}$ are sections of $L^{-1}$. It is covered by Example 1.1 via the mapping $L\to X\times\oplus_{1}^{m}{\mathbb{C}}$, defined by $(\sigma_{1},\ldots,\sigma_{m})$. Such metrics often occur in applications to complex algebraic geometry. ∎
We have that $s(E,h,h_{0})=s_{0}(E,h,h_{0})+s_{1}(E,h_{0},h)+\cdots$, where $s_{k}(E,h,h_{0})$ are closed $(k,k)$-currents of order 0. They are not necessarily positive unless both $h$ and $h_{0}$ are Griffiths negative. There are anyway well-defined multiplicities $\operatorname{mult}_{x}s_{k}(E,h,h_{0})$ at each point $x$, defined in the same way as Lelong numbers, see Section 4. Here is our first main theorem:
@@
Assume that $E\to X$ is a holomorphic vector bundle equipped with a qas metric $h$, and let $Z$ be the degeneracy locus of $h$. Let $h_{0}$ be a smooth reference metric, and let $s(E,h,h_{0})$ denote the associated Segre form.
(i) If $h_{\epsilon}$ is the metric defined by (1.1), then (1.2) holds.
@@
where $M^{E,h,h_{0}}$ is closed and has support on $Z$ and $s^{\prime}(E,h)={\bf 1}_{X\setminus Z}s(E,h,h_{0})$ is independent of $h_{0}$. If $h$ is smooth on $X\setminus Z$, then $s^{\prime}(E,h)$ coincides with the usual Segre form of $h$ there.
@@
(vi) For each integer $k\geq 0$ there is a decomposition
@@
where $S_{k}^{E,h}$ is the Lelong current of an analytic cycle of codimension $k$, and the set of $x$ such that $\operatorname{mult}_{x}s_{k}^{\prime}(E,h)$ or $\operatorname{mult}_{x}N_{k}^{E,h,h_{0}}$ is nonvanishing, is contained in an analytic set of codimension $\geq k+1$. The multiplicities of $s_{k}^{\prime}(E,h)$, $S_{k}^{E,h}$ and $N_{k}^{E,h,h_{0}}$ are non-negative integers that are independent of $h_{0}$.
(vii) If $h^{\prime}$ is a qas metric on $E\to X$ comparable to $h$, then $\operatorname{mult}_{x}M_{k}^{E,h,h_{0}}=\operatorname{mult}_{x}M_{k}^{E,h^{\prime},h_{0}}$ and $\operatorname{mult}_{x}s^{\prime}_{k}(E,h)=\operatorname{mult}_{x}s_{k}^{\prime}(E,h^{\prime})$ for $x\in X$ and $k=0,1,2,\ldots$. In particular, $S_{k}^{E,h}=S_{k}^{E,h^{\prime}}$, $k=0,1,2,\ldots$.
If the metric is $\mathbb{Q}$}-qas ($\mathbb{R}$}-qas), then the same statements hold, except that the multiplicities in part (vi) are rational (real).
@@
(ii) The properties (ii) to (iv) are (at least implicitly) proved in [18].
@@
(iv) If $M_{k}^{E,h,h_{0}}$ is a positive current, (1.7) is precisely the Siu decomposition. ∎
The definition of a Griffiths negative singular metric was introduced in [9], and its dual metric is by definition Griffiths positive. Although most attention in the literature is paid to positive metrics, we focus in this article on qas metrics which are slightly more general than negative metrics with integral analytic singularities. This leads to simpler formulations, in particular less signs. By considering duals of such metrics, one obtains corresponding statements concerning Griffiths positive metrics, cf. Remark 6.2.
Recall that in algebraic geometry, characteristic classes of vector bundles such as the Chern class, are defined as endomorphisms on Chow groups, i.e., the $\mathbb{Z}$}-module of analytic cycles ${\mathcal{Z}}(X)$ modulo rational equivalence.
Inspired by [5] we introduce a $\mathbb{Z}$}-module ${\mathcal{QZ}}(X)$ of currents of order 0 that contains the module ${\mathcal{Z}}(X)$ of analytic cycles as a submodule, if cycles are identified with their associated Lelong currents. In particular the $(0,0)$-current 11 is an element in ${\mathcal{QZ}}(X)$. Any element ${\mathcal{QZ}}(X)$ of a fixed dimension has a well-defined multiplicity (Lelong number) that is an integer, at each point. We also introduce certain quotient classes ${\mathcal{QB}}(X)$ that, contrary to the Chow classes, keep local information such as multiplicities; this fact is crucial in this paper.
A second main result is that the Segre form $s(E,h,h_{0})$ extends to an endomorphism ${\bf s}(E,h,h_{0})$ on ${\mathcal{QZ}}(X)$ such that ${\bf s}_{0}(E,h,h_{0})=I$, the identity, and ${\bf s}(E,h,h_{0})1=s(E,h,h_{0})$. An important fact is that ${\bf s}(E,h,h_{0})$ induces a well-defined endomorphism ${\bf s}(E,h)$ of ${\mathcal{QB}}(X)$ that is independent of $h_{0}$. There is an analogue of Theorem 1.3 for ${\bf s}(E,h,h_{0})$, see Theorem 6.7. Since ${\bf s}(E,h,h_{0})-I$ has positive degree,
@@
is a well-defined endomorphism such that
(1.9) ${\bf c}(E,h,h_{0}){\bf s}(E,h,h_{0})={\bf s}(E,h,h_{0}){\bf c}(E,h,h_{0})={I}.$
It is natural, cf. (2.4), to say that $c(E,h,h_{0}):={\bf c}(E,h,h_{0})1$ is the Chern form associated with $(E,h,h_{0})$. The same definition of $c(E,h,h_{0})$ appeared in [18] but expressed in a different way.
@@
Let $(E,h,h_{0})$ be as in Theorem 1.3. Properties (ii)-(vii) of Theorem 1.3 hold if $s(E,h,h_{0})$, $s^{\prime}(E,h)$, and $M^{E,h,h_{0}}$ are replaced by $c(E,h,h_{0})$, $c^{\prime}(E,h)$, and ${\bf 1}_{Z}s(E,h,h_{0})$, except that the multiplicities are not non-negative.
The analogue of (1.2) is not true for $c(E,h,h_{0})$, see Section 8.
As already mentioned, the starting point for this paper is [18] and the ideas there. The regularization (1.1) relies on results in [6, 10]. The arithmetic analysis of the Lelong numbers relies on ideas in [1], and the $\mathbb{Z}$}-modules ${\mathcal{QZ}}(X)$ and ${\mathcal{QB}}(X)$ that are modelled after the somewhat smaller modules ${\mathcal{GZ}}(X)$ and ${\mathcal{B}}(X)$ in [5].
@@ Acknowledgements
We are grateful to Håkan Samuelsson Kalm, Antonio Trusiani and David Witt Nyström for valuable discussions.
@@ 2. Preliminaries
Throughout this article $X$ is assumed to be a complex manifold. Given a holomorphic vector bundle $E\to X$ we let $p\colon{\mathbb{P}}(E)\to X$ denote the projectivization, so that the fiber above $x$ is the set of all lines [^2] There is another common convention in the literature, where the fiber instead is all hyperplanes through the origin. through the origin in $E_{x}$. Then the pullback $p^{*}E\to{\mathbb{P}}(E)$ has the tautological sub(line)bundle $L_{E}\to{\mathbb{P}}(E)$ whose fiber over $(x,[\alpha])\in{\mathbb{P}}(E)$ is the line in $E_{x}$ determined by $\alpha$.
If $E$ has a smooth Hermitian metric $h$, then $p^{*}E$ and hence $L_{E}$ inherit smooth Hermitian metrics. Notice that a local section $s$ of $L_{E}$ has the form $s=\tilde{s}(x,\alpha)\alpha$, where $\tilde{s}$ is a holomorphic function on $E\setminus 0_{E}$ which is 1-1-homogeneous in $\alpha$. Thus any local section can be identified with such an $\tilde{s}$. The induced norm is
@@
Therefore $dd^{c}\log|\alpha|_{h}^{2}$ is a well-defined form on ${\mathbb{P}}(E)$ and $c_{1}(L_{E},e^{\psi})=-dd^{c}\log|\alpha|^{2}_{h}$ is the first Chern form. Moreover,
$s(L_{E},e^{\psi})=\frac{1}{c(L_{E},e^{\psi})}=\frac{1}{1+c_{1}(L_{E},e^{\psi})}=\sum_{\ell=0}^{\infty}(dd^{c}\log|\alpha|_{h}^{2})^{\ell}.$
@@
and it is a smooth form since $p$ is a proper submersion. We write $s(E,h)=\sum s_{k}(E,h)$, where $s_{k}(E,h)$ has bidegree $(k,k)$. It turns out that $s_{0}(E,h)=1$ and thus s(E,h)1s(E,h)-1 has positive bidegree. We get the Chern form
@@
By [19], see also [14], this definition of $c(E,h)$ coincides with the definition as the sum of the elementary symmetric polynomials acting on ($i/2\pi$ times) the curvature tensor $\Theta(h)=D_{h}^{2}$ of $(E,h)$, where $D_{h}$ denotes the Chern connection associated with $h$. The Segre form is then defined as $s(E,h)=1/c(E,h)$.